OpenStudy (anonymous):

what is the value of x, y, and z for (5/4)x-(4/3)y+(5/4)z=6 , (5/16)x+(5/4)y-(4/16)z=-4 , (3/4)x-1y+(3/4)z=6

OpenStudy (anonymous):

\[(5/4)x-(4/3)y+(5/4)z=6 \times12\] \[15x -16y+15z=72\] \[(5/16)x+(5/4)y-(4/16)z=-4 \times 16\] \[5x+20y-4z=-64\] \[(3/4)x -y+(3/4)z=6 \times 4\] [3x-4y+3z=24 after that you can using subtitute or eliminate in order to get the value of x, y, z